OpenStudy (anonymous):

Use the distributive property to help you solve the equation 8(w-3) = -24

OpenStudy (saifoo.khan):

Hiiii.

OpenStudy (anonymous):

Hiiii!!!! Saif!

OpenStudy (saifoo.khan):

8w - 24 = -24

OpenStudy (anonymous):

8w-24=-24

OpenStudy (anonymous):

8(w-3) = -24 8w -24 = -24 add 24 to both sides 8w = 0 Hence w = 0

OpenStudy (anonymous):

so w = 0?

OpenStudy (saifoo.khan):

YuP!

OpenStudy (anonymous):

Thanks!